# Advanced Encryption Standard AI artificial intelligence

### v1

@@ 3.2 Step 2: Word-Level Reconstruction
To eliminate synthesis byproducts hindering analysis and enhance readability for human reverse engineers, we initially pre-processed the netlist.
@@
Figure 3: Netlist Preparation Overview.
For more efficient manual analysis and considering that some LUTs merely implement basic functions like AND or XOR, we substituted such LUTs with primitive gates, see 2(a). We also removed buffer LUTs as they add no functionality but sometimes hinder structural analysis. Furthermore, we removed gates and sub-circuits that compute the same function on identical inputs. Such duplicates are introduced by the synthesizer for optimization purposes, yet they can obfuscate shared control signals or data inputs.
MUXes in particular provide structure to the data path. On FPGAs, however, their functionality is implemented as part of more complex Boolean functions within LUTs. Hence, decomposing complex LUTs into small building blocks such as MUXes can aid manual and automated reverse engineering, e.g., by supporting structural analysis. To recover a nested MUX and better understand the data path (see Figure 5), we analyze the LUT’s Boolean function and search for inputs that behave like a select. When discovering a MUX in a LUT, we replace the LUT with a MUX and reconstruct the surrounding logic from primitive gates.
In total, we replaced 549 LUTs with primitive gates, removed 594 buffers, extracted 1619 MUXes and 1161 primitive gates, removed 353 duplicate gates, and removed 51 additional gates using other simplifications.
@@
Analyzing Maggie’s control path was required for comprehending how the circuit operates. Additionally, certain techniques operating on the data path require knowledge on which gates and nets belong to the control path. Unfortunately, the absence of FPGA-compatible reference implementations [34, 35, 36], the absence of public implementations altogether [37, 38, 39, 40], or general assumptions on strongly connected components that do not always hold for our case study [39, 38], prevented us from applying existing techniques from literature. In particular, the NETA toolset [41] is only available as binaries that cannot parse FPGA netlists. Moreover, a common assumption is that the state registers and transition logic of an FSMs form a SCC [39, 38]. For Maggie, we found 55 SCCs with all but seven comprising three gates or less. Vast parts of the netlist form a single SCC of more than 5000 gates. Therefore, SCCs are not a useful metric in our case study.
Given the intricacies of this issue, we ourselves could not find a workable solution within the scope of this work and hence decided to manually investigate the control path of Maggie instead. Here, we faced significant challenges due to the FSMs being interwoven to an extent that often prevented even manual separation, as discussed in [40]. While FSM separation was not always feasible, we still classified FFs as part of the control path by checking whether they feed into control pins such as MUX select or FF enable. After assigning a FF to the control path, we also examined its predecessors and successors. These structural characteristics are merely clues that a gate might belong to the control path. They often still required manual inspection and human intuition, hence preventing full automation.
@@
Investigating Maggie’s SPI, UART, and I2S interface implementations (see Figure 1) supports word-level structure recovery as these interfaces often comprise the first or last register of a data path. Furthermore, we must reverse engineer these interfaces to understand the parallel and serial conversion of incoming and outgoing data, which is precondition for recovering bit orders and data types. Through manual reverse engineering, we identified 615 gates belonging to five interfaces.
@@
In the iPhone netlist, we observed the data path comprising mostly combinational logic implemented as LUTs and carry gates as well as sequential logic implemented using FFs. Some data-path structures are implemented in a consistent, recurring way, e.g., adders, counters, and selected comparators. Lattice iCE40 FPGAs provide carry gates that are commonly connected to form carry chains, often used in combination with LUTs to implement arithmetic operations, see Figure 4. A reverse engineer can leverage these recurring pattern to identify word-level structures like adders or counters in a gate-level netlist.
Figure 4: Structure of a $3$-bit counter with reset constructed from a carry chain on a Lattice iCE40 FPGA.
To recover word-level arithmetic operations, we first locate carry chains in the netlist by scanning for connected carry gates. We then assemble arithmetic structure candidates by adding surrounding logic to the carry chains. Since a structure may be influenced by control inputs, we construct structural candidate variants to account for different control behavior. Operand membership as well as the order of input and output signals are derived from the structure and connectivity of a candidate’s gates. For example, the order of the input signals is reconstructed by identifying the carry gate of the chain that a signal is connected to. Finally, we check each candidate against predefined satisfiability modulo theories (SMT) models, e.g., $A\pm B$ for adders or $A\pm n$ for counters incrementing by a constant $n$.
@@
The synthesizer realizes word-level registers from the HDL design as individual FFs that lack any indication of which register they belong to. To this end, DANA [42] can be used to recover these word-level registers from a gate-level netlist using structural metrics such as shared predecessors and successors as well as common control signals. However, as DANA is optimized for data-path analysis, control FFs may not be recognized as such and instead be wrongfully merged into data-path registers. This impairs the detection of other registers due to DANA’s iterative nature. Hence, we extended DANA to include known registers identified by other analysis techniques for finding, e.g., communication interfaces, FSMs states, and counters. DANA does not alter these registers, but utilizes them to identify registers further up or down the data path.
@@
In the data path, MUXes can be used to switch between word-level data sources, such as global inputs, memory, or arithmetic structures. We implemented a method to group MUXes that have been decomposed from LUTs as part of our pre-processing, see Figure 3. A MUX selects an input to be forwarded to the output based on a select input S. Multiple MUXes sharing a common select typically form a word-level MUX structure. Detecting such structures allows for the separation of word-level data paths, see Figure 5. However, grouping by select can yield oversized structures as the same select signal may influence two or more distinct data paths. Hence, we split these groups according to preceding or successive word-level structures. For example, if a 32-bit MUX is succeeded by two 16-bit registers, the MUX is split into two 16-bit MUXes. As other preceding or successive MUXes themselves are considered for splitting, the process is repeated iteratively until no further changes can be observed.
Figure 5: The recovery of word-level MUXes allows for the analysis of separate independent data paths.
@@
We propagated the bit order from structures with known bit order along the data path to structures with unknown bit order. For every unordered pin group, our algorithm ($i$) derives bit-order candidates from predecessors and successors and ($ii$) tries to establish consensus between them. Candidates that have an inherent conflict are discarded, e.g., when different pins of a pin group fetched the same index from the same origin. We omit control logic during analysis, as it does not feature a natural bit order and degrades results.
@@
Arithmetic
35 692
Register
118 1342
Multiplexers
160 1217
Control Logic
- 631
TABLE I: Results of our word-level reconstruction
In addition to the presented methods, we manually grouped 688 gates into data-path modules that did not feature MUXes. In total, we assigned 4570 out of 5496 gates (83%) to structures using manual and automated techniques. Furthermore, we assigned a bit order to 605 (71%) out of 843 pin groups from word-level structures that we did not previously classify as control logic. While we could not verify against a ground truth because of the black-box nature of our case study, these results enabled correctly extracting the implemented algorithm, thereby underlining their value.
3.3 Step 3: Algorithmic Recovery
Having recovered word-level structures in the previous step, we now extract an algorithmic description of what is implemented by the iPhone netlist. This entails further analyzing the control path, as we could not draw boundaries between individual control structures so far.
@@ 3.3 Step 3: Algorithmic Recovery
Instead of further attempting to unravel the semantics of the interwoven control logic, we decided to use dynamic reverse engineering techniques to observe the actual behavior of the circuit at runtime. However, such dynamic analysis often requires expensive equipment and yields only incomplete insights into the execution state [43, 20]. Hence, we leveraged a more workable approach: netlist simulation using I/O signals captured during operation. We further used this approach to ($i$) observe the flow of real-world data through the netlist, assisting in algorithmic recovery, and ($ii$) to verify correctness of the recovered netlist and the later extracted algorithm implemented on Maggie by comparing simulation results against recorded outputs.
@@
We captured I/O signals by placing a custom breakout board between Maggie and the iPhone PCB, see Figure 7. By feeding the captured I/O signals to a simulator, we could then analyze the state of the circuit at any point in time. This reduced complexity of the circuit’s theoretical state space by focusing only on the states that are actually reached. To this end, we extended the netlist analysis framework HAL [24] with simulation capabilities, see Appendix A. This approach allowed for, e.g., analysis of the data flow over time and correct traversal of states for complex FSMs that otherwise exhausted functional analysis. Given the previously reconstructed word-level structures and their bit orders, we could now follow the flow of multi-bit data through the design and observe which operations were executed at what point in time. This completely alleviated us from the need to statically analyze the control path.
@@
By tracing inputs to Maggie through the netlist in simulation, we found that almost all data-path operations are performed by four interconnected DSPs. On Lattice iCE40 FPGAs, DSPs feature four 16-bit data inputs and a 32-bit output [44]. On Maggie, all DSPs are configured as Multiply-Accumulate (MAC) units performing 16-bit multiplication followed by 32-bit accumulation. Most DSP inputs are preceded by LUTs implementing MUXes to allow for dynamically changing data sources, see Figure 8. Each DSP and its preceding MUXes are controlled by their own FSM. Looking at the simulation, we found that every DSP computes a recurring sequence of operations depending on dynamically changing control and data inputs.
Figure 8: Simplified DSP circuit.
@@
3.4 Step 4: High-Level Sense-Making
Based on the recovered algorithm, we proceed with a semantic analysis that is essential for meaningful interaction with Apple’s IP. Generating a Python script implementing the algorithm proved crucial, since comprehending the algorithm required domain knowledge that was beyond our FPGA reverse engineering team. This way, we could share the Python script containing the reconstructed algorithm with a signal processing expert without requiring them to have any knowledge of the reverse engineering process.
@@ 3.4 Step 4: High-Level Sense-Making
To generate the LRA control signals, Maggie operates on two feedback signals from the LRA, a reference signal previously loaded into BRAM, and values of intermediate signals at previous sampling instants. From the algorithmic representation, we identified several classical biquad filter stages implemented in a direct form I structure [45]. Upon closer inspection, we found many typical DSP building blocks, e.g., sub-sampling and sample-and-hold blocks for sample rate conversion, bit shifts and truncations for fixed-point arithmetic, saturation stages to prevent overflows, and filter implementations. Furthermore, we recognized even subtle details such as the use of fraction saving [46]. Putting it all together, we found three main processing stages: $(i)$ input signal conditioning and application of calibration, $(ii)$ a state observer to track the actuator dynamics, and $(iii)$ the closed-loop controller to minimize errors between the actual and the desired LRA movement. For details on the recovered DSP operations, see Appendix B.
Figure 9: Overview of Maggie’s signal processing.
4 Deriving Generalized Techniques
@@
Our case study demonstrated that a semi-automated approach for FPGA reverse engineering in the context of IP theft entails considerable effort. In line with RQ2, we now investigate whether (and to what extent) the techniques from our case study can be across different FPGA architectures and implementations. We provide open-source implementations of the tools presented in this section as plugins to HAL [24] upon publication of this paper.
We evaluated the techniques discussed in this section on seven benchmark designs synthesized for Xilinx 7-series and Lattice iCE40 FPGAs, see Table II for details on the benchmarks and Table III for evaluation results. Differences in resource utilization between the Xilinx and Lattice benchmarks, e.g., for fft64, can be the result of the Xilinx synthesizer using DSPs where the Lattice synthesizer resorts to carry chains. Furthermore, Xilinx FPGAs leverage 6-input LUTs while Lattice FPGAs use 4-input ones.
4.1 Netlist Pre-Processing
@@
So far, our pre-processing steps were limited to the 4-input LUTs found on Lattice iCE40 FPGAs. While some techniques such as the detection of duplicate gates could easily be extended to cover arbitrary combinational gates, this is not true for our LUT decomposition. Here, our approach works reasonably well for 4-input LUTs, but does not scale to six inputs. Hence, for generalization we re-synthesize the combinational subgraphs between sequential gates using open-source EDA tools instead of analyzing each LUT’s Boolean function. To this end, we constrain the synthesizer to a custom gate library comprising only a the desired combinational gates such as AND, OR, XOR, and MUX of various sizes. Compared to decomposing the combinational parts of the netlist into a pure and-or-invert (AOI) graph, this approach allows us to delegate multiplexer (MUX) detection to the synthesizer by incentivizing the selection of MUX gates over alternatives by artificially reducing their size and optimizing for area. This way, we can even identify MUXes within complex LUT configurations. By nature, this re-synthesis approach is independent of the underlying FPGA architecture and target implementation.
@@
The pre-processing is a best effort approach and lacks any ground truth to compare against. It is, however, a precondition for subsequent techniques such as word-level MUX detection and arithmetic structure identification. The success of these techniques serves as a proxy to gauge the effectiveness of our pre-processing approach.
4.2 Arithmetic Structures
Figure 10: Module Identification and Classification Overview.
In our case study we identified arithmetic operations by building candidate sets of gates and verifying their function using an SMT solver. However, utilizing an SMT solver to compare against functional models from a library of known operations requires knowledge of the order of input and output signals as well as their operand membership. Furthermore, identifying any existing control signals presents additional challenges. When analyzing Maggie, we inferred operands of arithmetic operations from structural properties such as the pins of carry gates, e.g., operand $A$ of an addition $A+B$ would often be connected to input I0 of a Lattice iCE40 CARRY gate. The bit order of arithmetic operations was deduced from the structure of the underlying carry chains. This approach is similar to an idea proposed by Narayanan et al. [47]. However, such structural properties always depend on the particular implementation, FPGA architecture, surrounding logic, and synthesizer optimizations, hence calling for a more generic solution.

### v2

@@ 3.2. Step 2: Word-Level Reconstruction
We initially pre-processed the netlist to eliminate synthesis byproducts hindering analysis and enhance readability for human reverse engineers.
@@
Figure 3. Netlist Preparation Overview.
For more efficient manual analysis and considering that some LUTs merely implement basic functions like AND or XOR, we substituted such LUTs with primitive gates, see 3(a). We also removed buffer LUTs as they add no functionality but sometimes hinder structural analysis. Furthermore, we removed gates and sub-circuits that compute the same function on identical inputs. The synthesizer introduces such duplicates for optimization purposes, yet they can obfuscate shared control signals or data inputs.
MUXes in particular provide structure to the data path. On FPGAs, however, their functionality is implemented as part of more complex Boolean functions within LUTs. Hence, decomposing complex LUTs into small building blocks such as MUXes can aid manual and automated reverse engineering, e.g., by supporting structural analysis. To recover nested MUXes and better understand the data path (see Figure 5), we analyzed each LUT’s Boolean function and searched for inputs that behave like a select. When discovering a MUX in a LUT, we replaced the LUT with a MUX and reconstruct the surrounding logic from primitive gates. In total, we replaced 549 LUTs with primitive gates, removed 594 buffers, extracted 1619 MUXes and 1161 primitive gates, removed 353 duplicate gates, and removed 51 additional gates using other simplifications.
@@
Analyzing Maggie’s control path was required for comprehending how the circuit operates. Additionally, certain techniques operating on the data path require knowledge of which gates and nets belong to the control path. Unfortunately, the absence of FPGA-compatible reference implementations (Meade et al., 2018a, 2016b; Geist et al., 2020), the absence of public implementations altogether (Brunner et al., 2019; Fyrbiak et al., 2018; Shi et al., 2010; Brunner et al., 2022), or general assumptions on strongly connected components that do not always hold for our case study (Shi et al., 2010; Fyrbiak et al., 2018), prevented us from applying existing techniques from literature. In particular, the NETA toolset (Meade et al., 2018b) is only available as binaries that cannot parse FPGA netlists. Moreover, a common assumption is that the state registers and transition logic of an FSMs form an SCC (Shi et al., 2010; Fyrbiak et al., 2018). For Maggie, we found 55 SCCs with all but seven comprising three gates or less. Vast parts of the netlist form a single SCC of more than 5000 gates. Therefore, SCCs are not useful in our case study.
Given the intricacies of this issue, we could not find a workable solution within the scope of this work and hence decided to manually investigate the control path of Maggie instead. Here, we faced significant challenges due to the FSMs being interwoven to an extent that often prevented even manual separation, as discussed in (Brunner et al., 2022). While FSM separation was not always feasible, we still classified FFs as part of the control path by checking whether they feed into control pins such as MUX select or FF enable. After assigning a FF to the control path, we examined its predecessors and successors. These structural characteristics are merely clues that a gate might belong to the control path. They often required manual inspection and human intuition, hence preventing full automation.
@@
Investigating Maggie’s SPI, UART, and I2S interface implementations (see Figure 1) supports word-level structure recovery as these interfaces often comprise the first or last register of a data path. Furthermore, we must reverse engineer these interfaces to understand the parallel and serial conversion of incoming and outgoing data, which is a precondition for recovering bit orders and data types. Through manual reverse engineering, we identified 615 gates belonging to five interfaces.
@@
In the iPhone netlist, we observed the data path comprising mostly combinational logic implemented as LUTs and carry gates as well as sequential logic implemented using FFs. Some data-path structures are implemented in a consistent, recurring way, e.g., adders, counters, and selected comparators. Lattice iCE40 FPGAs provide carry gates that are commonly connected to form carry chains, often used in combination with LUTs to implement arithmetic operations, see Figure 4. A reverse engineer can leverage these recurring patterns to identify word-level structures like adders or counters in a gate-level netlist.
Figure 4. Structure of a $3$-bit counter with reset constructed from a carry chain on a Lattice iCE40 FPGA.
To recover word-level arithmetic operations, we first located carry chains in the netlist by scanning for connected carry gates. We then assembled arithmetic structure candidates by adding surrounding logic to the carry chains. Since a structure may be influenced by control inputs, we constructed structural candidate variants to account for different control behaviors. Operand membership and the order of input and output signals were derived from the structure and connectivity of a candidate’s gates. For example, the order of inputs was reconstructed by identifying the carry gate of the chain that a signal is connected to. Finally, we checked each candidate against predefined satisfiability modulo theories (SMT) models, e.g., $A\pm B$ for adders or $A\pm n$ for counters incrementing by a constant $n$.
@@
The synthesizer realizes word-level registers from the HDL design as individual FFs that lack any indication of which register they belong to. To this end, DANA (Albartus et al., 2020) can be used to recover these word-level registers from a gate-level netlist using structural metrics such as shared predecessors and successors as well as common control signals. However, as DANA is optimized for data-path analysis, control FFs may not be recognized as such and instead be wrongfully merged into data-path registers. This impairs the detection of other registers due to DANA’s iterative nature. Hence, we extended DANA to include known registers identified by other analysis techniques for finding, e.g., communication interfaces, FSMs states, and counters. DANA does not alter these registers but utilizes them to identify registers further up or down the data path.
@@
In the data path, MUXes can be used to switch between word-level data sources, such as global inputs, memory, or arithmetic structures. We implemented a method to group MUXes that have been decomposed from LUTs as part of our pre-processing, see Figure 3. A MUX selects an input to be forwarded to the output based on a select input S. Multiple MUXes sharing a common select typically form a word-level MUX structure. Detecting such structures allows for the separation of word-level data paths, see Figure 5. However, grouping by select can yield oversized structures as the same select signal may influence two or more distinct data paths. Hence, we split these groups according to preceding or successive word-level structures. For example, if a 32-bit MUX was succeeded by two 16-bit registers, the MUX was split into two 16-bit MUXes. As other preceding or successive MUXes themselves were considered for splitting, the process was repeated iteratively until no further changes could be observed.
Figure 5. The recovery of word-level MUXes allows for the analysis of separate independent data paths.
@@
We propagated the bit order from structures with known bit order along the data path to structures with unknown bit order. For every unordered pin group, our algorithm ($i$) derives bit-order candidates from predecessors and successors and ($ii$) tries to establish consensus between them. Candidates with inherent conflict are discarded, e.g., when different pins of a pin group fetched the same index from the same origin. We omit control logic during analysis, as it does not feature a natural bit order and degrades results.
@@
Table 1 depicts the results for our automated techniques. In addition, we manually grouped 688 gates into data-path modules that did not feature MUXes and identified 631 gates belonging to control logic. In total, we assigned 4570 out of 5496 gates (83%) to structures using manual and automated techniques. Furthermore, we assigned a bit order to 605 (71%) out of 843 pin groups from word-level structures that were not classified as control logic. While we could not verify against a ground truth as of the black-box nature of our case study, these results enabled correctly extracting the implemented algorithm, thereby underlining their value.
@@
Arithmetic 35 692
Register 118 1342
Multiplexers 160 1217
Table 1. Results of our word-level reconstruction.
3.3. Step 3: Algorithmic Recovery
Having previously recovered word-level structures, we extracted an algorithmic description of what the iPhone netlist implements. This entails further analyzing the control path, as we could not draw boundaries between individual control structures so far.
@@ 3.3. Step 3: Algorithmic Recovery
Instead of further attempting to unravel the semantics of the interwoven control logic, we decided to use dynamic reverse engineering techniques to observe the actual behavior of the circuit at runtime. However, such dynamic analysis often requires expensive equipment and yields only incomplete insights into the execution state (Nedospasov et al., 2012; Tajik et al., 2017). Hence, we leveraged a more workable approach: netlist simulation using I/O signals captured during operation. We further used this approach to ($i$) observe the flow of real-world data through the netlist, assisting in algorithmic recovery, and ($ii$) to verify correctness of the recovered netlist and the later extracted algorithm implemented on Maggie by comparing simulation results against recorded outputs.
@@
We captured I/O signals by placing a custom breakout board between Maggie and the iPhone PCB, see Figure 7. By feeding the captured I/O signals to a simulator, we could then analyze the state of the circuit at any point in time. This reduced complexity of the circuit’s theoretical state space by focusing only on the states that are actually reached. To this end, we extended the netlist analysis framework HAL (HAL, 2018) with simulation capabilities, see Appendix A. This approach allowed for, e.g., analysis of the data flow over time and correct traversal of states for complex FSMs that otherwise exhausted functional analysis. Given the previously reconstructed word-level structures and their bit orders, we could now follow the flow of multi-bit data through the design and observe which operations were executed at what point in time. This completely alleviated us from the need to statically analyze the control path.
@@
By tracing inputs to Maggie through the netlist in simulation, we found that almost all data-path operations are performed by four connected DSPs. On Lattice iCE40 FPGAs, DSPs feature four 16-bit data inputs and a 32-bit output (Semiconductor, 2016). On Maggie, all DSPs are configured as Multiply-Accumulate (MAC) units performing 16-bit multiplication and 32-bit accumulation. Most DSP inputs are preceded by LUTs implementing MUXes to allow for dynamically changing data sources, see Figure 8. Each DSP and its MUXes are controlled by an FSM. Looking at the simulation, we found that every DSP computes a recurring sequence of operations depending on dynamically changing control and data inputs.
Figure 8. Simplified DSP circuit.
@@
3.4. Step 4: High-Level Sense-Making
Based on the recovered algorithm, we proceeded with a semantic analysis that is essential for meaningful interaction with Apple’s IP. Generating a Python script implementing the algorithm proved crucial, since comprehending the algorithm required domain knowledge that was beyond our FPGA reverse engineering team. This way, we could share the Python script containing the reconstructed algorithm with a signal processing expert without requiring them to have any knowledge of the reverse engineering process.
@@ 3.4. Step 4: High-Level Sense-Making
To generate the LRA control signals, Maggie operates on two feedback signals from the LRA, a reference signal previously loaded into BRAM, and values of intermediate signals at previous sampling instants. From the algorithmic representation, we identified several classical biquad filter stages implemented in a direct form I structure (Oppenheim and Schafer, 2014). Upon closer inspection, we found many typical DSP building blocks, e.g., sub-sampling and sample-and-hold blocks for sample rate conversion, bit shifts and truncations for fixed-point arithmetic, saturation stages to prevent overflows, and filter implementations. Furthermore, we recognized even subtle details such as the use of fraction saving (Yates and Lyons, 2008). Putting it all together, we found three main processing stages: $(i)$ input signal conditioning and application of calibration, $(ii)$ a state observer to track the actuator dynamics, and $(iii)$ the closed-loop controller to minimize errors between the actual and the desired LRA movement. For details on the recovered DSP operations, see Appendix B.
Figure 9. Overview of Maggie’s signal processing.
4. Deriving Generalized Techniques
@@
Our case study demonstrated that a semi-automated approach for FPGA reverse engineering in the context of IP theft entails considerable effort. In line with RQ2, we now investigate whether (and to what extent) the learnings from our case study can be generalized across different FPGA architectures. Upon publication, we will provide open-source implementations of our generalized tools.
We evaluated the techniques discussed in this section on six benchmark designs synthesized for Xilinx 7-series and Lattice iCE40 FPGAs, see Table 2 for evaluation results and Appendix C for details on the benchmarks. Furthermore, we conducted an end-to-end white-box case study on one signal processing design implemented for Xilinx 7-series FPGAs to showcase the effectiveness of our generalized techniques in concert.
4.1. Netlist Pre-Processing
@@
So far, our pre-processing steps were limited to the 4-input LUTs found on Lattice iCE40 FPGAs. While some techniques such as the detection of duplicate gates could easily be extended to cover arbitrary combinational gates, this is not true for our LUT decomposition. Here, our approach works reasonably well for 4-input LUTs but does not scale to six inputs. Hence, for generalization, we re-synthesize the combinational subgraphs between sequential gates using open-source EDA tools instead of analyzing each LUT’s Boolean function. To this end, we constrain the synthesizer to a custom gate library comprising only the desired combinational gates such as AND, OR, XOR, and MUX of various sizes. Compared to decomposing the combinational parts of the netlist into a pure and-or-invert (AOI) graph, this approach allows us to delegate multiplexer (MUX) detection to the synthesizer by incentivizing the selection of MUX gates over alternatives by artificially reducing their size and optimizing for area. This way, we can even identify MUXes within complex LUT configurations. By nature, this re-synthesis approach is independent of the underlying FPGA architecture and target implementation.
@@
We verified the correctness of our pre-processing using SMT solving. However, pre-processing is a best-effort approach and lacks a ground truth to compare against. It is essential for subsequent techniques such as word-level MUX detection and arithmetic structure identification. The success of these techniques serves as a proxy to gauge the effectiveness of our pre-processing approach.
4.2. Arithmetic Structures
Figure 10. Module Identification and Classification Overview.
In our case study, we identified arithmetic operations by building candidate sets of gates and verifying their function using an SMT solver. However, utilizing an SMT solver to compare against functional models from a library of known operations requires knowledge of the order of input and output signals as well as their operand membership. Furthermore, identifying any existing control signals presents additional challenges. When analyzing Maggie, we inferred operands of arithmetic operations from structural properties such as the pins of carry gates, e.g., operand $A$ of an addition $A+B$ would often be connected to input I0 of a Lattice iCE40 CARRY gate. The bit order of arithmetic operations was deduced from the structure of the underlying carry chains. This approach is similar to an idea proposed by Narayanan et al. (Narayanan et al., 2023). However, such structural properties always depend on the particular implementation, FPGA architecture, surrounding logic, and synthesizer optimizations, hence calling for a more generic solution.